A wire has resistance of 3.1 ohm at 30°C and 4.5 ohm at 100°C .The temperature coefficient of resistance to the wire is ...?

Grade:11

2 Answers

Sanju
6 years ago
R1 = 3.1 ohm at t = 300C
R2 = 4.5 ohm at t = 1000C
We have, R = R0 (1+αt)
So, R= R0 (1+α * 30)
R= R0 (1+α * 100)
R/ R= [(1+30α) / (1+100α)]
3.1 /4.5 = [(1+30α) / (1+100α)] 
Solving the above expression gives:
Temp. coefficient, α = 0.0085 0C-1
ankit singh
3 years ago
R= R0 (1+α * 30)
R= R0 (1+α * 100)
R/ R= [(1+30α) / (1+100α)]
3.1 /4.5 = [(1+30α) / (1+100α)]
solve for coefficient